OpenStudy (askme12345):

The same meter stick is used as a rigid pendulum with no weight. What is its period?

OpenStudy (askme12345):

@isaiah.feynman

OpenStudy (anonymous):

t = 1/f f = 1/2PIE *SQRT (K/m)

OpenStudy (anonymous):

Insufficient information

OpenStudy (isaiah.feynman):

No @iambatman there is more than enough info. The period will be the same if there were no weight. The period of a pendulum is independent of the weight.